# Taming Rectified Flow for Inversion and Editing

### v2

@@ Abstract
Rectified-flow-based diffusion transformers like FLUX and OpenSora have demonstrated outstanding performance in the field of image and video generation. Despite their robust generative capabilities, these models often struggle with inversion inaccuracies, which could further limit their effectiveness in downstream tasks such as image and video editing. To address this issue, we propose RF-Solver, a novel training-free sampler that effectively enhances inversion precision by mitigating the errors in the ODE-solving process of rectified flow. Specifically, we derive the exact formulation of the rectified flow ODE and apply the high-order Taylor expansion to estimate its nonlinear components, significantly enhancing the precision of ODE solutions at each timestep. Building upon RF-Solver, we further propose RF-Edit, a general feature-sharing-based framework for image and video editing. By incorporating self-attention features from the inversion process into the editing process, RF-Edit effectively preserves the structural information of the source image or video while achieving high-quality editing results. Our approach is compatible with any pre-trained rectified-flow-based models for image and video tasks, requiring no additional training or optimization. Extensive experiments across generation, inversion, and editing tasks in both image and video modalities demonstrate the superiority and versatility of our method. The source code is available at this URL.
{strip}[Uncaptioned image]
Figure 1: RF-Solver for downstream tasks in image and video. We propose RF-Solver to solve the rectified flow ODE with reduced error, thereby enhancing both sampling quality and inversion-reconstruction accuracy for rectified-flow-based generative models [2, 3]. Furthermore, we propose RF-Edit, which utilizes the RF-Solver for image and video editing tasks. Our methods demonstrate impressive performance across generation, inversion, and editing tasks in both image and video modalities.
@@ 1 Introduction
Recent advancements of generation methods based on Rectified Flow (RF) [42, 35, 67] have demonstrated exceptional performance in synthesizing high-quality images and videos. Different from traditional approaches represented by Stable Diffusion [22, 56], these methods leverage the Diffusion Transformer [51, 78, 73, 63] architecture and implement a straight-line motion system to produce the desired data distribution. With these effective designs, FLUX [2] and OpenSora [3] have respectively emerged as one of the state-of-the-art (SOTA) methods in the field of Text-to-Image (T2I) and Text-to-Video (T2V) generation.
@@
Despite the remarkable success in the fundamental T2I and T2V generation tasks, few studies have explored the performance of RF-based models on various downstream tasks such as inversion-reconstruction [60, 47, 17, 71] and editing [20, 45]. When directly applying the vanilla RF for inversion, we observe that it fails to faithfully reconstruct the image or video from the source. Examples are shown in Fig. 1 Task 1 and Task 2 (the third row). For image inversion, the positions of objects (e.g. the cake) and the appearance of individuals (e.g. the child) in the reconstructed image significantly diverge from the source image. The performance of video inversion is even worse, with noticeable distortions present in the reconstructed video. The inaccuracies of inversion and reconstruction would severely constrain the performance of RF models on other inversion-based downstream tasks such as image editing [20, 64, 48, 13, 29] and video editing [41, 33, 15, 59].
@@
In this work, we investigate the aforementioned problem by delving into the inversion and reconstruction process of the RF. Specifically, we track the latent at each intermediate timestep during inversion and reconstruction, calculating the Mean Square Error (MSE) between them at corresponding timesteps. We observe that significant errors are introduced at each timestep throughout the whole reconstruction process, and their accumulation ultimately results in a considerably deviated output (the red curve in Fig. 2). Based on the definition and inference process of RF [40, 42], we identify that these errors stem from the Ordinary Differential Equation (ODE) solving process [79, 66, 23, 52]. Specifically, the essence of the inversion and generation process for RF is to derive the solution of RF ODE [42]. Since this ODE includes terms involving complex neural networks, the solution can only be coarsely approximated by a sampler. However, the experiment in Fig. 2 indicates that the sampler adopted in existing models [2, 3] lacks sufficient precision for the inversion task, causing notable errors to accumulate at each timestep, finally leading to unsatisfactory reconstruction results.
@@
Based on the analysis, we aim to improve inversion accuracy by introducing a more effective sampler, which is more general and fundamental than designing a specific inversion method. To this end, we propose RF-Solver. Specifically, we note that the exact formulation of the RF ODE solution can be directly derived using the variation of constants method. For the nonlinear component of this solution (i.e., the integral of the neural network), we utilize Taylor expansion for estimation. By employing higher-order Taylor expansion, the ODE can be solved with reduced error, thereby enhancing the performance of RF models. RF-Solver is a generic sampler that can be seamlessly integrated into any rectified flow model without additional training or optimization. Experimental results demonstrate that RF-Solver not only significantly enhances the accuracy of inversion and reconstruction (the green curve in Fig. 2), but also improves performance on fundamental tasks such as T2I generation.
Building upon this, we propose RF-Edit to leverage RF-Solver in editing tasks. Real-world image and video editing require the model to make precise modifications to a source image/video while maintaining its overall structure unchanged, presenting greater challenges than reconstruction. In this scenario, it is inadequate to solely rely on the inverted noises as prior knowledge for editing, which could lead to edited results being excessively influenced by the target prompt, diverging significantly from the original source [20, 64]. Addressing this problem, RF-Edit stores the $\mathcal{V}$ (value) feature in the self-attention layers at several timesteps during inversion. These features are used to replace the corresponding features in the denoising process. Practically, we design two specific sub-modules for RF-Edit, respectively leveraging the DiT structure of FLUX [2] and OpenSora [3] as the backbones for image and video editing. With the effective design of RF-Edit, it demonstrates superior performance in both image and video domains, outperforming various SOTA methods.
Our core contributions are summarized as follows:
  • We propose RF-Solver, a training-free sampler that significantly reduces errors in the inversion and reconstruction processes of rectified-flow models.
  • We present RF-Edit, which leverages RF-Solver for image and video editing. RF-Edit effectively preserves the structural integrity of the source image/video while delivering high-quality results.
  • Extensive experiments across a range of tasks demonstrate the efficacy and versatility of our method.
@@ 2.1 Inversion
Inversion maps the real visual data, i.e. image and video, to representations in noise space, which is the reverse process of generation. The representative method, DDIM inversion [60, 61], adds predicted noise recursively at each forward step. Many efforts [14, 65, 47, 57, 46, 27, 46, 44, 19] have been made to mitigate the discretization error in DDIM inversion. Despite the effectiveness of inversion in diffusion models, the exploration of inversion in SOTA rectified flow models like FLUX and OpenSora is limited. RF-prior [76] uses the score distillation to invert the image while it requires many optimizing steps. More recently, [58] introduces an additional vector field conditioned on the source image to improve the inversion. However, the error from the original vector field of rectified flow still persists, which would limit the performance of such method on various downstream tasks. In contrast, we aim to directly mitigate the error from the original vector field in this work.
@@ 2.2 Image and Video Editing
Training-free methods for image and video editing [25, 62] have gained increasing popularity for their efficiency and effectiveness. Existing image editing methods focus on prompt refinement [55, 69], attention-sharing mechanism [20, 50, 18, 64], mask guidance [4, 10, 24, 37, 19], and noise initialization [5, 77]. Video editing introduces additional complexities in maintaining temporal consistency, making it a more challenging task. Existing video editing methods focus on attention injection [53, 70, 41], motion guidance [9, 16, 75, 68], latent manipulation [81, 74, 30, 8], and canonical representation [7, 36, 49, 31]. To date, the editing performance of RF-based diffusion transformers has remained largely under-explored. Although [58] employs FLUX [2] for image editing, its performance is limited to simple tasks such as stylization and face editing while often failing to effectively maintain the structural information of source images. Moreover, currently there is no research exploring the video editing capabilities of RF-based models.
@@ 3.1 Preliminaries
Rectified Flow (RF) [43] facilitates the transition between the real data distribution $\pi_{0}$ and Gaussian Noises distribution $\pi_{1}$ along a straight path. This is achieved by learning a forward-simulating system defined by the ODE: $d\boldsymbol{Z}_{t}=v(\boldsymbol{Z}_{t},t)dt,t\in[0,1]$, which maps $\boldsymbol{Z}_{1}\in\pi_{1}$ to $\boldsymbol{Z}_{0}\in\pi_{0}$.
In practice, the velocity field $v$ is parameterized by a neural network $\boldsymbol{v}_{\theta}$. During training, given empirical observations of two distributions $\boldsymbol{X}_{0}\sim\pi_{0}$, $\boldsymbol{X}_{1}\sim\pi_{1}$ and $t\in[0,1]$, the forward process (i.e. adding noise) of rectified flow is defined by a simple linear combination: $\boldsymbol{X}_{t}=t\boldsymbol{X}_{1}+(1-t)\boldsymbol{X}_{0}$. The differential form of the equation is given by: $d\boldsymbol{X}_{t}=(\boldsymbol{X}_{1}-\boldsymbol{X}_{0})dt$. Consequently, the training process optimizes the network by solving the least squares regression problem, which fits the $\boldsymbol{v}_{\theta}$ with $(\boldsymbol{X}_{1}-\boldsymbol{X}_{0})$:
@@
The RF model can generate high-quality images in much fewer timesteps compared to DDPM [22], owing to the nearly linear transition trajectory established during training. With these effective designs, RF model illustrates great potential in the field of T2I and T2V generation [2, 3].
@@ 3.2 RF-Solver
The vanilla RF sampler demonstrates strong performance in image and video generation. However, when applied to inversion and reconstruction tasks, we observe significant error accumulation at each timestep. This results in reconstructions that diverge notably from the original image (see Fig. 2). This severely limits the performance of RF models in various inversion-based downstream tasks [20, 66]. Delving into this problem, we identify that the errors stem from the process of estimating the approximate solution for the rectified flow ODE [67, 40], which is formulated by Eq. 2 in existing methods [2, 3]. Consequently, obtaining more precise solutions for the ODE would effectively mitigate these errors, leading to improved reconstruction quality.
@@
where $\boldsymbol{v}^{(k)}_{\theta}(\boldsymbol{Z}_{t_{i}},t_{i})=\frac{\text{d}^{k}\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}},t_{i})}{\text{d}t^{k}}$, denoting the $k$-order derivative of $\boldsymbol{v}_{\theta}$ and $\mathcal{O}$ denotes higher-order infinitesimals. Substituting Eq. 4 into the integral term yields:
@@
Substituting Eq. 6 and Sec. 3.2 into Eq. 3, we derive the $n$-th order solution of Rectified flow ODE:
@@
where $h_{i}:=t_{i-1}-t_{i}$. Sec. 3.2 indicates that to estimate $\boldsymbol{Z}_{t_{i-1}}$, we need to obtain the $k$-th order derivatives $\{\boldsymbol{v}^{(k)}_{\theta}(\boldsymbol{Z}_{t_{i}},{t_{i}})\}$ for $k\in\{0,\cdots,n-1\}$.
When $n=1$, the formula reduces to the standard rectified flow (i.e., Eq. 2). In our experiments, we find that setting $n=2$ effectively mitigates the errors, yielding:
@@
Figure 3: RF-Edit pipelines for image editing and video editing. We design two sub-modules for applying RF-Edit to (a). Image editing with FLUX [2] and (b). Video editing with OpenSora [3]. Note that for FLUX, there are multiple Double Blocks, followed by multiple Single Blocks. For OpenSora, there are multiple OpenSora DiT blocks. For simplicity, only one block of each type is depicted in the figure.
Note that $\boldsymbol{v}_{\theta}^{(1)}$ in Sec. 3.2 is the first-order derivative of the network prediction term $\boldsymbol{v}_{\theta}$, which cannot be analytically derived due to the complex architecture of the neural network. To estimate this term, we first obtain the network prediction $\hat{\boldsymbol{v}}_{t_{i}}$ at the timestep $t_{i}$, i.e., $\hat{\boldsymbol{v}}_{t_{i}}=\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}},{t_{i}})$. Then we step forward a small timestep $\Delta t=\frac{1}{2}(t_{i-1}-t_{i})$, and update the latents to obtain $\boldsymbol{Z}_{{t_{i}}+\Delta t}=\boldsymbol{Z}_{t_{i}}+\Delta t\cdot\hat{\boldsymbol{v}}_{t_{i}}$.
Subsequently, we calculate an additional prediction of the network at the timestep $t_{i}+\Delta t$, i.e., $\hat{\boldsymbol{v}}_{t_{i}+\Delta t}=\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}+\Delta t},{t_{i}+\Delta t})$. With $\hat{\boldsymbol{v}}_{t_{i}}$ and $\hat{\boldsymbol{v}}_{t_{i}+\Delta t}$, the first-order derivative of ${\boldsymbol{v}}_{\theta}$ at the timestep $t_{i}$ can be estimated as: $\boldsymbol{v}_{\theta}^{(1)}(\boldsymbol{Z}_{t_{i}},t_{i})=\frac{\hat{\boldsymbol{v}}_{t_{i}+\Delta t}-\hat{\boldsymbol{v}}_{t_{i}}}{\Delta t}$. Substituting this formulation into Sec. 3.2 results in the practical implementation of the RF-Solver algorithm. The complete sampling process for RF-Solver is presented in Algorithm 1.
Algorithm 1 Sampling process of RF-Solver
Input:
  $\boldsymbol{v}_{\theta}$triangleright Velocity function
  $t=[t_{N},\ldots,t_{0}]$triangleright Time steps
For $i=N$ to $1$ do
  $\Delta t_{i}\leftarrow\frac{1}{2}(t_{i-1}-t_{i})$
Output: $\boldsymbol{Z}_{0}$
Obtaining the sampling form of RF-Solver, we further derive its inversion form. Inversion maps data back into noise, which reverses the sampling process. Following previous methods for DDIM inversion [60, 12], the ODE process can be directly reversed in the limit of small steps. Based on this assumption, the inversion process of RF-Solver (Sec. 3.2) can be directly derived as:
@@
where $\widetilde{\boldsymbol{Z}}_{t_{i}}$ and $\widetilde{\boldsymbol{Z}}_{t_{i+1}}$ denotes the latents during inversion. Through the high order expansion, the error of the ODE solution in each timestep is reduced from $\mathcal{O}\big{(}(h_{i})^{2}\big{)}$ to $\mathcal{O}\big{(}(h_{i})^{3}\big{)}$, leading to improved performance, particularly in inversion and reconstruction (see Fig. 2). Beyond inversion and reconstruction, RF-Solver can also be applied to any RF-based model (such as FLUX [2] and OpenSora [3]) for other tasks such as sampling and editing, enhancing performance without requiring additional training.

### v3

@@ Abstract
Rectified-flow-based diffusion transformers like FLUX and OpenSora have demonstrated outstanding performance in the field of image and video generation. Despite their robust generative capabilities, these models often struggle with inversion inaccuracies, which could further limit their effectiveness in downstream tasks such as image and video editing. To address this issue, we propose RF-Solver, a novel training-free sampler that effectively enhances inversion precision by mitigating the errors in the ODE-solving process of rectified flow. Specifically, we derive the exact formulation of the rectified flow ODE and apply the high-order Taylor expansion to estimate its nonlinear components, significantly enhancing the precision of ODE solutions at each timestep. Building upon RF-Solver, we further propose RF-Edit, a general feature-sharing-based framework for image and video editing. By incorporating self-attention features from the inversion process into the editing process, RF-Edit effectively preserves the structural information of the source image or video while achieving high-quality editing results. Our approach is compatible with any pre-trained rectified-flow-based models for image and video tasks, requiring no additional training or optimization. Extensive experiments across generation, inversion, and editing tasks in both image and video modalities demonstrate the superiority and versatility of our method. Code is available at this URL .
Machine Learning, ICML
[…]
Junfu Pujevinpu@tencent.com
{strip}
Figure 1: RF-Solver for downstream tasks in image and video. We propose RF-Solver to solve the rectified flow ODE with reduced error, thereby enhancing both sampling quality and inversion-reconstruction accuracy for rectified-flow-based generative models (Black-Forest-Labs, 2024; Zheng et al., 2024). Furthermore, we propose RF-Edit, which utilizes the RF-Solver for editing. Our methods demonstrate impressive performance across generation, inversion, and editing tasks in both image and video modalities.
@@ 1 Introduction
Recent advancements of generation methods based on Rectified Flow (RF) (Liu et al., 2022a; Lee et al., 2024; Wang et al., 2024b) have demonstrated exceptional performance in synthesizing high-quality images and videos. Different from traditional approaches represented by Stable Diffusion (Ho et al., 2020; Rombach et al., 2022), these methods leverage the Diffusion Transformer (Peebles & Xie, 2023; Yang et al., 2024c; Xie et al., 2024; Tang et al., 2024) architecture and implement a straight-line motion system to produce the desired data distribution. With these effective designs, FLUX (Black-Forest-Labs, 2024) and OpenSora (Zheng et al., 2024) have respectively emerged as one of the state-of-the-art (SOTA) methods in the field of Text-to-Image (T2I) and Text-to-Video (T2V) generation.
@@
Despite the remarkable success in the fundamental T2I and T2V generation tasks, few studies have explored the performance of RF-based models on various downstream tasks such as inversion-reconstruction (Song et al., 2021a; Mokady et al., 2023; Guo et al., 2024; Wang et al., 2023c) and editing (Hertz et al., 2022; Meng et al., 2022). When directly applying the vanilla RF for inversion, we observe that it fails to faithfully reconstruct the image or video from the source. Examples are shown in Figure 1 Task 1 and Task 2 (the third row). For image inversion, the positions of objects (e.g., the cake) and the appearance of individuals (e.g., the child) in the reconstructed image significantly diverge from the source image. The performance of video inversion is even worse, with noticeable distortions present in the reconstructed video. The inaccuracies of inversion and reconstruction would severely constrain the performance of RF models on other inversion-based downstream tasks such as image editing (Hertz et al., 2022; Tumanyan et al., 2023; Nguyen et al., 2024; Duan et al., 2024; Ju et al., 2024) and video editing (Liu et al., 2024; Ku et al., 2024; Fan et al., 2024; Shin et al., 2024).
@@
In this work, we investigate the aforementioned problem by delving into the inversion and reconstruction process of the RF. Specifically, we track the latent at each intermediate timestep during inversion and reconstruction, calculating the Mean Square Error (MSE) between them at corresponding timesteps. We observe that significant errors are introduced at each timestep throughout the whole reconstruction process, and their accumulation ultimately results in a considerably deviated output (the red curve in Figure 2). Based on the definition and inference process of RF (Liu, 2022; Liu et al., 2022a), we identify that these errors stem from the Ordinary Differential Equation (ODE) solving process (Zhang et al., 2024; Wang et al., 2024a; Hong et al., 2024; Peng et al., 2024). Specifically, the essence of the inversion and generation process for RF is to derive the solution of RF ODE (Liu et al., 2022a). Since this ODE includes terms involving complex neural networks, the solution can only be coarsely approximated by a sampler. However, the experiment in Figure 2 indicates that the sampler adopted in existing models (Black-Forest-Labs, 2024; Zheng et al., 2024) lacks sufficient precision for the inversion task, causing notable errors to accumulate at each timestep, finally leading to unsatisfactory reconstruction results.
@@
Based on the analysis, we aim to improve inversion accuracy by introducing a more effective sampler, which is more general and fundamental than designing a specific inversion method. To this end, we propose RF-Solver. Specifically, we note that the exact formulation of the RF ODE solution can be directly derived using the variation of constants method. For the nonlinear component of this solution (i.e.,, the integral of the neural network), we utilize Taylor expansion for estimation. By employing higher-order Taylor expansion, the ODE can be solved with reduced error, thereby enhancing the performance of RF models. RF-Solver is a generic sampler that can be seamlessly integrated into any rectified flow model without additional training or optimization. Experimental results demonstrate that RF-Solver not only significantly enhances the accuracy of inversion and reconstruction (the green curve in Figure 2), but also improves performance on fundamental tasks such as T2I generation.
Building upon this, we propose RF-Edit to leverage RF-Solver in editing tasks. Real-world image and video editing require the model to make precise modifications to a source image/video while maintaining its overall structure unchanged, presenting greater challenges than reconstruction. In this scenario, it is inadequate to solely rely on the inverted noises as prior knowledge for editing, which could lead to edited results being excessively influenced by the target prompt, diverging significantly from the original source (Hertz et al., 2022; Tumanyan et al., 2023). Addressing this problem, RF-Edit stores the $\mathcal{V}$ (value) feature in the self-attention layers at several timesteps during inversion. These features are used to replace the corresponding features in the denoising process. Practically, we design two specific sub-modules for RF-Edit, respectively leveraging the DiT structure of FLUX (Black-Forest-Labs, 2024) and OpenSora (Zheng et al., 2024) as the backbones for image and video editing. With the effective design of RF-Edit, it demonstrates superior performance in both image and video domains, outperforming various SOTA methods.
@@ 2.1 Inversion
Inversion maps the real visual data, i.e., image and video, to representations in noise space, which is the reverse process of generation. The representative method, DDIM inversion (Song et al., 2021a, b), adds predicted noise recursively at each forward step. Many efforts (Elarabawy et al., 2022; Wallace et al., 2023; Mokady et al., 2023; Rout et al., 2024a; Miyake et al., 2023; Lu et al., 2022) have been made to mitigate the discretization error in DDIM inversion. Despite the effectiveness of inversion in diffusion models, the exploration of inversion in SOTA rectified flow models like FLUX and OpenSora is limited. RF-prior (Yang et al., 2024b) uses the score distillation to invert the image while it requires many optimizing steps. More recently, (Rout et al., 2024b) introduces an additional vector field conditioned on the source image to improve the inversion. However, the error from the original vector field of rectified flow still persists, which would limit the performance of such method on various downstream tasks. In contrast, we aim to directly mitigate the error from the original vector field in this work.
@@ 2.2 Image and Video Editing
Training-free methods for image and video editing (Huang et al., 2024a; Sun et al., 2024) have gained increasing popularity for their efficiency and effectiveness. Existing image editing methods focus on prompt refinement (Ravi et al., 2023; Wang et al., 2023a), attention-sharing mechanism (Hertz et al., 2022; Parmar et al., 2023; Cao et al., 2023; Tumanyan et al., 2023), mask guidance (Avrahami et al., 2023; Couairon et al., 2022; Huang et al., 2023), and noise initialization (Brack et al., 2023; Yang et al., 2023b). Video editing introduces additional complexities in maintaining temporal consistency, making it a more challenging task. Existing video editing methods focus on attention injection (Qi et al., 2023; Wang et al., 2023b; Liu et al., 2024), motion guidance (Cong et al., 2023; Geyer et al., 2023; Yang et al., 2024a; Wang et al., 2024c), latent manipulation (Zhang et al., 2023; Yang et al., 2023a; Kara et al., 2024; Chen et al., 2023), and canonical representation (Chai et al., 2023; Lee et al., 2023; Ouyang et al., 2024; Kasten et al., 2021). To date, the editing performance of RF-based diffusion transformers has remained largely under-explored. Although (Rout et al., 2024b) employs FLUX (Black-Forest-Labs, 2024) for image editing, its performance is limited to simple tasks such as stylization and face editing while often failing to effectively maintain the structural information of source images. Moreover, currently there is no research exploring the video editing capabilities of RF-based models.
@@ 3.1 Preliminaries
Rectified Flow (RF) (Liu et al., 2022b) facilitates the transition between the real data distribution $\pi_{0}$ and Gaussian Noises distribution $\pi_{1}$ along a straight path. This is achieved by learning a forward-simulating system defined by the ODE: $d\boldsymbol{Z}_{t}=v(\boldsymbol{Z}_{t},t)dt,t\in[0,1]$, which maps $\boldsymbol{Z}_{1}\in\pi_{1}$ to $\boldsymbol{Z}_{0}\in\pi_{0}$.
In practice, the velocity field $v$ is parameterized by a neural network $\boldsymbol{v}_{\theta}$. During training, given empirical observations of two distributions $\boldsymbol{X}_{0}\sim\pi_{0}$, $\boldsymbol{X}_{1}\sim\pi_{1}$ and $t\in[0,1]$, the forward process (i.e., adding noise) of rectified flow is defined by a simple linear combination: $\boldsymbol{X}_{t}=t\boldsymbol{X}_{1}+(1-t)\boldsymbol{X}_{0}$. The differential form of the equation is given by: $d\boldsymbol{X}_{t}=(\boldsymbol{X}_{1}-\boldsymbol{X}_{0})dt$. Consequently, the training process optimizes the network by solving the least squares regression problem, which fits the $\boldsymbol{v}_{\theta}$ with $(\boldsymbol{X}_{1}-\boldsymbol{X}_{0})$:
@@
The RF model can generate high-quality images in much fewer timesteps compared to DDPM (Ho et al., 2020), owing to the nearly linear transition trajectory established during training. With these effective designs, RF model illustrates great potential in the field of T2I and T2V generation (Black-Forest-Labs, 2024; Zheng et al., 2024).
@@ 3.2 RF-Solver
The vanilla RF sampler demonstrates strong performance in image and video generation. However, when applied to inversion and reconstruction tasks, we observe significant error accumulation at each timestep. This results in reconstructions that diverge notably from the original image (see Figure 2). This severely limits the performance of RF models in various inversion-based downstream tasks (Hertz et al., 2022; Wang et al., 2024a). Delving into this problem, we identify that the errors stem from the process of estimating the approximate solution for the rectified flow ODE (Wang et al., 2024b; Liu, 2022), which is formulated by Equation 2 in existing methods (Black-Forest-Labs, 2024; Zheng et al., 2024). Consequently, obtaining more precise solutions for the ODE would effectively mitigate these errors, leading to improved performance.
@@
where $\boldsymbol{v}^{(k)}_{\theta}(\boldsymbol{Z}_{t_{i}},t_{i})=\frac{\text{d}^{k}\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}},t_{i})}{\text{d}t^{k}}$, denoting the $k$-order derivative of $\boldsymbol{v}_{\theta}$ and $\mathcal{O}$ denotes higher-order infinitesimals. Substituting Equation 4 into the integral term yields:
@@
Substituting Equation 6 and Section 3.2 into Equation 3, we derive the $n$-th order solution of Rectified flow ODE:
@@
where $h_{i}:=t_{i-1}-t_{i}$. Section 3.2 indicates that to estimate $\boldsymbol{Z}_{t_{i-1}}$, we need to obtain the $k$-th order derivatives $\{\boldsymbol{v}^{(k)}_{\theta}(\boldsymbol{Z}_{t_{i}},{t_{i}})\}$ for $k\in\{0,\cdots,n-1\}$.
Considering $n=1$, the formula reduces to the standard rectified flow (i.e.,, Equation 2). In our experiments, we find that setting $n=2$ effectively mitigates the errors, yielding:
@@
Figure 3: RF-Edit pipelines for image editing and video editing. We design two sub-modules for applying RF-Edit to (a). Image editing with FLUX (Black-Forest-Labs, 2024) and (b). Video editing with OpenSora (Zheng et al., 2024). Note that for FLUX, there are multiple Double Blocks, followed by multiple Single Blocks. For OpenSora, there are multiple OpenSora DiT blocks. For simplicity, only one block of each type is depicted in the figure.
Note that $\boldsymbol{v}_{\theta}^{(1)}$ in Section 3.2 is the first-order derivative of the network prediction term $\boldsymbol{v}_{\theta}$, which cannot be analytically derived due to the complex architecture of the neural network. To estimate this term, we first obtain the network prediction $\hat{\boldsymbol{v}}_{t_{i}}$ at the timestep $t_{i}$, i.e., $\hat{\boldsymbol{v}}_{t_{i}}=\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}},{t_{i}})$. Then we step forward a small timestep $\Delta t$ (which is set to $0.01$ in experiments), and update the latents to obtain $\boldsymbol{Z}_{{t_{i}}+\Delta t}=\boldsymbol{Z}_{t_{i}}+\Delta t\cdot\hat{\boldsymbol{v}}_{t_{i}}$. Subsequently, we calculate an additional prediction of the network at the timestep $t_{i}+\Delta t$, i.e., $\hat{\boldsymbol{v}}_{t_{i}+\Delta t}=\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}+\Delta t},{t_{i}+\Delta t})$. With $\hat{\boldsymbol{v}}_{t_{i}}$ and $\hat{\boldsymbol{v}}_{t_{i}+\Delta t}$, the first-order derivative of ${\boldsymbol{v}}_{\theta}$ at the timestep $t_{i}$ can be estimated as: $\boldsymbol{v}_{\theta}^{(1)}(\boldsymbol{Z}_{t_{i}},t_{i})=\frac{\hat{\boldsymbol{v}}_{t_{i}+\Delta t}-\hat{\boldsymbol{v}}_{t_{i}}}{\Delta t}$. Substituting this formulation into Section 3.2 results in the practical implementation of the RF-Solver algorithm. The complete sampling process for RF-Solver is presented in Algorithm 1.
Obtaining the sampling form of RF-Solver, we further derive its inversion form. Inversion maps data back into noise, which reverses the sampling process. Following previous methods for DDIM inversion (Song et al., 2021a; Dhariwal & Nichol, 2021), the ODE process can be directly reversed in the limit of small steps. Based on this assumption, the inversion process of RF-Solver can be derived as:
@@
where $\widetilde{\boldsymbol{Z}}_{t_{i}}$ and $\widetilde{\boldsymbol{Z}}_{t_{i+1}}$ denotes the latents during inversion. Through the high order expansion, the error of the ODE solution in each timestep is reduced from $\mathcal{O}\big{(}(h_{i})^{2}\big{)}$ to $\mathcal{O}\big{(}(h_{i})^{3}\big{)}$, significantly facilitating inversion and reconstruction process (see Figure 2). Beyond this, RF-Solver can also be applied to any RF-based model (such as FLUX (Black-Forest-Labs, 2024) and OpenSora (Zheng et al., 2024)) for other tasks such as sampling and editing, enhancing performance without requiring additional training.